# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.2.2 The (relatively) high-error case
    for a sufficiently large constant $C^{\prime}_{1}>0$, which holds whenever $k$ is larger than some appropriate absolute constant. Instantiating Lemma 2.23 with $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ above yields a strong $(k_{1}=0.8n_{1},k_{2}=0.8n_{2},\varepsilon_{1}+\varepsilon_{2})$-block-source extractor $\mathsf{BExt}\colon\{0,1\}^{n_{1}}\times\{0,1\}^{n_{2}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m_{1}}$.
    Since $X^{\prime}$ is $3\varepsilon^{\prime}$-close to an $(n_{1},n_{2},0.8)$-block source, we conclude that
    $Y_{\mathsf{BExt}}\circ\mathsf{BExt}(X^{\prime},Y_{\mathsf{BExt}})\approx_{3\varepsilon^{\prime}+\varepsilon_{1}+\varepsilon_{2}}U_{d_{2}+m_{1}}.$ (9)
We define the output of our final strong extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m_{1}}$ to be $\mathsf{BExt}(X^{\prime},Y_{\mathsf{BExt}})$. Since $\varepsilon_{1},\varepsilon_{2}\leq\varepsilon^{\prime}$, Equation 9 implies that
$Y_{\mathsf{Cond}}\circ Y_{\mathsf{BExt}}\circ\mathsf{Ext}(X,Y_{\mathsf{Cond}}\circ Y_{\mathsf{BExt}})\approx_{5\varepsilon^{\prime}}U_{d+m_{1}}.$
This means that $\mathsf{Ext}$ is a strong $(k,\varepsilon^{\prime}+5\varepsilon^{\prime}=\varepsilon)$-seeded extractor with seed length $d=|Y_{\mathsf{Cond}}|+|Y_{\mathsf{BExt}}|=O(\log(n/\varepsilon))$ and output length $m_{1}\geq c_{1}k_{1}\geq c^{\prime}_{1}k$ for an absolute constant $c^{\prime}_{1}>0$, where one of the $\varepsilon^{\prime}$ terms in the error comes from fixing the seed in the condensing step of Item 1.
@@
Finally, we analyze the time complexity of $\mathsf{Ext}$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then Item 1 runs in time $\widetilde{O}(n)$. In Item 3, $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ are both computable in time $\widetilde{O}(n)$ under this lower bound on $k$, and thus so is $\mathsf{BExt}$. We conclude that $\mathsf{Ext}$ runs in time $\widetilde{O}(n)$. Otherwise, if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then Item 1 runs in time $\widetilde{O}(n)$ after a preprocessing step, and $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ in Item 3 run in time $\widetilde{O}(n)$ after a preprocessing step. Therefore, overall, $\mathsf{Ext}$ runs in time $\widetilde{O}(n)$ after a preprocessing step. ∎

### v2

@@ 4.2.2 The (relatively) high-error case
    for a sufficiently large constant $C^{\prime}_{1}>0$, which holds whenever $k$ is larger than some appropriate absolute constant. Instantiating lemma˜2.22 with $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ above yields a strong $(k_{1}=0.8n_{1},k_{2}=0.8n_{2},\varepsilon_{1}+\varepsilon_{2})$-block-source extractor $\mathsf{BExt}\colon\{0,1\}^{n_{1}}\times\{0,1\}^{n_{2}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m_{1}}$.
    Since $X^{\prime}$ is $3\varepsilon^{\prime}$-close to an exact $(n_{1},n_{2},0.8)$-block source, we conclude that
    $\mathopen{}\mathclose{{\left(Y_{\mathsf{BExt}},\mathsf{BExt}(X^{\prime},Y_{\mathsf{BExt}})}}\right)\approx_{3\varepsilon^{\prime}+\varepsilon_{1}+\varepsilon_{2}}U_{d_{2}+m_{1}}.$ (20)
We define the output of our final strong extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m_{1}}$ to be $\mathsf{BExt}(X^{\prime},Y_{\mathsf{BExt}})$. Since $\varepsilon_{1}=\varepsilon_{2}=\varepsilon^{\prime}$, eq.˜20 implies that
$\mathopen{}\mathclose{{\left(Y_{\mathsf{Cond}},Y_{\mathsf{BExt}},\mathsf{Ext}(X,Y_{\mathsf{Cond}},Y_{\mathsf{BExt}})}}\right)\approx_{5\varepsilon^{\prime}}U_{d+m_{1}}.$
This means that $\mathsf{Ext}$ is a strong $(k,\varepsilon^{\prime}+5\varepsilon^{\prime}=\varepsilon)$-seeded extractor with seed length $d=|Y_{\mathsf{Cond}}|+|Y_{\mathsf{BExt}}|=O(\log(n/\varepsilon))$ and output length $m_{1}\geq c_{1}k_{1}\geq c^{\prime}_{1}k$ for an absolute constant $c^{\prime}_{1}>0$, where one of the $\varepsilon^{\prime}$ terms in the error comes from fixing the seed in the condensing step of item˜1.
@@
Finally, we analyze the time complexity of $\mathsf{Ext}$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon\geq 2^{-Cn^{0.1}}$, then item˜1 runs in time $\widetilde{O}(n)$. In item˜3, $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ are both computable in time $\widetilde{O}(n)$ under these lower bounds on $k$ and $\varepsilon$, and thus so is $\mathsf{BExt}$. We conclude that $\mathsf{Ext}$ runs in time $\widetilde{O}(n)$.
Otherwise, if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then item˜1 runs in time $\widetilde{O}(n)$ after a preprocessing step. And if $\varepsilon<2^{-Cn^{0.1}}$ then $\mathsf{Ext}_{1}$ and $\mathsf{Ext}_{2}$ in item˜3 run in time $\widetilde{O}(n)$ after a preprocessing step. Therefore, overall, $\mathsf{Ext}$ runs in time $\widetilde{O}(n)$ after a preprocessing step in this case. ∎
5 A Faster Instantiation of Trevisan’s Extractor
Definition 5.1 (weak design).
$\sum_{j<i}2^{|S_{i}\cap S_{j}|}\leq\rho(m-1).$
$\mathsf{Tre}(x,y)=(\bar{x}|_{y_{S_{1}}},\dots,\bar{x}|_{y_{S_{m}}}),$ (21)
where we denote $\bar{x}=\mathcal{C}(x)$ and interpret each length-$\log\bar{n}$ bit-string $y_{S_{i}}$ as a location in $\bar{n}]$. For the runtime analysis, it will be important to recall that $\delta$ is set to be $\frac{\varepsilon}{cm}$ for some universal constant $c$.
Theorem 5.2.
Trevisan’s extractor of eq.˜21, set to extract $m$ bits with any error $\varepsilon>0$, is computable in time $\widetilde{O}(n+m\log(1/\varepsilon))$.
Proof.
Looking at eq.˜21, note that we only need to compute $m$ coordinates of $\mathcal{C}(x)$. To compute those $m$ coordinates, $y_{S_{1}},\ldots,y_{S_{m}}$, we first need to compute the weak design itself. Note that this can be seen as a preprocessing step, since it only depends on the parameters of the extractor, and not on $x$ or $y$. We will use the following result.
Claim 5.3 ([FYE+25], Section A.5).
Once we have our preprocessing step, we are left with computing the code. By corollary˜3.3, we can choose $\bar{n}$ so that $n/\bar{n}=\delta^{c}$ for some universal constant $c$, and so $\bar{n}=n\cdot\operatorname{poly}(m,1/\varepsilon)$ and $\ell=\log\bar{n}=O(\log(n/\varepsilon))$. Generating the design can then be done in time $\widetilde{O}(m\log(n/\varepsilon))$. Now, corollary˜3.3 tells us that any $m$ bits of $\mathcal{C}(x)$ can be computed in time
$\widetilde{O}(n)+m\log(1/\delta)\cdot\operatorname{polylog}(n)=\widetilde{O}(n+m\log(1/\varepsilon)).$
On a RAM in the logarithmic cost model, we can use the variant of $\mathcal{C}$ that uses Spielman’s code as a base code (see remark˜3.4) and get a runtime of $O(n)+m\log(1/\varepsilon)\cdot\operatorname{polylog}(n)$. This gives a truly linear time construction whenever $m$ is at most $\frac{n}{\log(1/\varepsilon)\operatorname{polylog}(n)}$. ∎
Corollary 5.4.